From the following dissociation question:
Al2(SO4)3 ----> 2Al^3 + 3SO4^2-

What will be the concentration of the CATION present if 25.5g aluminum sulfate is dissolved in 1.00L of water?
1
Expert's answer
2020-06-02T14:08:21-0400

Solution:

The dissociation equation:

Al2(SO4)3 ---> 2Al3+ + 3SO42- 

Al3+ is the cation.

According to the equation: n(Al2(SO4)3) = n(Al3+)/2


Moles of Al2(SO4)3 = n(Al2(SO4)3) = Mass of Al2(SO4)3 / Molar mass of Al2(SO4)3

The molar mass of Al2(SO4)3 is 342.15 g/mol

n(Al2(SO4)3) = (25.5 g) / (342.15 g/mol) = 0.07453 mol


n(Al3+) = 2 × n(Al2(SO4)3) = 2 × 0.07453 mol = 0.14906 mol

Concentration of Al3+ = CM(Al3+) = Moles of Al3+ / Volume of solution

CM(Al3+) = (0.14906 mol) / (1.00 L) = 0.14906 mol/L = 0.149 mol/L

The concentration of Al3+ is 0.149 mol/L


Answer: The concentraion of Al3+ is 0.149 mol/L.